# Making Reasoning Matter: Measuring and Improving Faithfulness of Chain-of-Thought Reasoning

### v2

@@ Abstract
Large language models (LLMs) have been shown to perform better when asked to reason step-by-step before answering a question. However, it is unclear to what degree the model’s final answer is faithful to the stated reasoning steps. In this paper, we perform a causal mediation analysis on twelve LLMs to examine how intermediate reasoning steps generated by the LLM influence the final outcome and find that LLMs do not reliably use their intermediate reasoning steps when generating an answer. To address this issue, we introduce Frodo, a framework to tailor small-sized LMs to generate correct reasoning steps and robustly reason over these steps. Frodo consists of an inference module that learns to generate correct reasoning steps using an implicit causal reward function, and a reasoning module that learns to faithfully reason over these intermediate inferences using a counterfactual and causal preference objective. Our experiments show that Frodo significantly outperforms four competitive baselines. Furthermore, Frodo improves the robustness and generalization ability of the reasoning LM, yielding higher performance on out-of-distribution test sets. Finally, we find that Frodo’s rationales are more faithful to its final answer predictions than standard supervised fine-tuning.
@@
Debjit Paul, Robert West, Antoine Bosselut, Boi Faltings EPFL {firstname.lastname}@epfl.ch https://debjitpaul.github.io/reasoningmatter
@@ 1 Introduction
Figure 1: An example of our proposed causal analysis to measure the faithfulness of the final output with respect to the CoT generated by the model. We perturbed CoT rationales and studied the causal impact on the model’s behaviour.
Reasoning implicitly involves two steps: identifying the rules and facts (inference chains) necessary to reach a conclusion and then robustly using them to reach said conclusion (Levesque, 1986). Our paper studies whether LLMs reliably use inference chains to arrive at a conclusion. [^1] Please note that throughout our paper, when we mention reasoning faithfulness, we mean how reliably the model uses their reasoning steps to arrive at a correct answer. In standard CoT, LLMs can generate plausible explanations with the final answer not necessarily guaranteed to follow the reasoning chain or imply a causal relation between the reasoning chain and the model’s outcome (Lyu et al., 2023). Most recent efforts have either focused on the performance of LLMs on various reasoning tasks or their faithfulness in CoT generation, ignoring the sequential relationship between CoT and the final answer (Huang and Chang, 2023; Lanham et al., 2023).
@@
We apply this framework and study the causal impact of CoT rationales on the behaviour of twelve different state-of-the-art LLMs on three different kinds of complex reasoning tasks (mathematical, commonsense, and causal understanding). We observe a large variation across tasks and models in how strongly reasoning traces causally affect the model’s prediction. In particular, we find that instruction-tuned models (GPT-3.5-Instruct, Brown et al., 2020b; Mistral-Instruct-7B, Jiang et al., 2023) have a stronger causal effect on the final answer when conditioned on the reasoning trace than models trained with RLHF (e.g., ChatGPT; Llama-2-7B-Chat, Touvron et al., 2023). Similar to Turpin et al. (2023), when we intervene in the reasoning problem, we observe that ChatGPT and GPT-$3.5$-Instruct are inconsistent at generating plausible reasoning chains. Finally, we find GPT-$4$ (Achiam et al., 2023) only changes its answer $30$% of the time when conditioned on perturbed counterfactual reasoning chains. In Figure 1, we see one example where GPT-4 does not faithfully change its final answer when provided with intervened counterfactual CoT. These results indicate two issues: (i) LLMs can generate an unfaithful and implausible reasoning chains, and (ii) LLMs are inconsistent when reasoning over their own generated reasoning traces.
@@
We evaluate Frodo on four reasoning tasks (Quarel, StrategyQA, OpenBookQA, QASC) using multiple model backbones of different scales, and demonstrate that Frodo achieves an absolute accuracy improvement of $2\%\sim 3\%$ over baselines trained with standard supervised fine-tuning or chain-of-thought distilled methods. By intervening in reasoning chains, we assess the model’s robustness. We evaluate the models’ ability to change their answers when provided with counterfactual reasoning chains. Notably, Frodo exhibited significant (+$4.5\%$) improvement in robustness. Finally, Frodo generalizes better to out-of-distribution test sets, showing a +$2.6\%$ performance improvement over standard supervised fine-tuning. We will make our code and data available once the paper is accepted.
@@ 2 Reasoning Chain as a Mediator
The causal graph is a probabilistic graphical model used to describe how variables interact, expressed by a directed acyclic graph consisting of the sets of nodes ($N$) denoting the variables and the edges ($E$) between the nodes denoting the causality (Pearl, 1998). As shown in Fig. 2(a), if $Y$ is a descendent of $X$, then $X$ is a potential cause of $Y$, and $Y$ is the effect.
@@
It is a method of causal inference which studies the change in a dependent variable following an intervention or treatment of an independent variable (Pearl, 2001). Causal mediation analysis aims to decompose the total effect of the independent variable ($X$) on the dependent variable ($Y$) into two components: the direct effect and the indirect effect (Pearl, 2001). In this work, we view the reasoning process as a causal graph, framing the input (reasoning problem) $X$ and the output $Y$ as random variables and the reasoning steps as mediator variable $R$. We use mediation analysis to interpret the role of reasoning steps as mediators between model inputs and model outputs. Please note that there are two mediator variables, model parameters $M$ and reasoning steps $R$ (see Fig. 2(b)). However, in this work, we are interested in understanding the indirect effect of $R$ on the output $Y$; hence, we assume model parameters $M$ to be fixed. Let $X_{0}$ denote the initial reasoning problem, $R_{0}$ as the reasoning chain given $X_{0}$ and $Y_{00}$ denote the potential outcome if the treatment variable and mediator variable are $X_{0}$ and $R_{0}$, respectively. Whereas $Y_{01}$ denotes the possible outcome when treatment is set to $X_{0}$ and $R_{1}$ is the reasoning chain for an intervened reasoning problem (see Fig.2(d)).
@@
Figure 2: Causal graph for natural language reasoning, where we model $P(Y|do(x))$. $X_{0}$ = original reasoning problem, $X_{1}$ = intervened reasoning problem, $R_{0}$ = reasoning steps for $X_{0}$ reasoning problem, $R_{1}$ = reasoning steps for $X_{1}$, $Y$ = output, and $M$ = model parameters
@@
Direct Effect (DE) measures how much an intervention $X$ changes an outcome variable $Y$ directly, without passing through a hypothesized mediator $R$. Based on Fig 2(c), we define the direct effect of $X$=$X_{0}$ on $Y$ as $DE=P(Y|X_{0},R_{0})-P(Y|X_{1},R_{0})$ which can be seen as the comparisons between two potential outcomes of $X$ given two different treatments, i.e., $X$ = $X_{0}$ and $X_{1}$. It is computed by applying the intervention $X$ but holding $R$ fixed to its original value ($R_{0}$).
@@
Indirect Effect measures how much the intervention $X$ changes $Y$ indirectly through $R$. We define the indirect effect as $IE=P(Y|X_{0},R_{0})-P(Y|X_{0},R_{1})$. It is computed by setting $R$ to its value under the intervention $X$ while keeping everything else to its original value.
@@
More concretely, a high direct effect means that $X$ has higher causality on $Y$, i.e., the model puts more emphasis on the reasoning problem than the reasoning steps. In contrast, a high indirect effect means the model puts more emphasis on the reasoning steps than the problem input.
@@
Table 1: An example from StrategyQA dataset, where $X_{1}$, $R_{0}$ and $R_{1}$ are generated by GPT-4.
Reasoning Intervention.
Following Pearl (2001), we conduct counterfactual reasoning to measure the causal effect of a treatment variable on a response variable. We first perform targeted interventions on the input text $X$ and measure their effect on the reasoning outcome $Y$ (direct effect). Further, we also perform interventions on the mediator $R$ and measure their effect on $Y$ (indirect effect). We perform the following steps to automatically generate an intervention on $X$ and $R$.
Step 1: Intervention Data Generation. We use a large language model (GPT-4) to automatically generate an alternative treatment variable $X_{1}$ [^2] Please note that we have performed different interventions specific to each reasoning task (details in App. A.4).. The input to $M$ includes instruction and few-shot examples, taking the format shown in Table 11. LLMs can be sensitive to instructions and few-shot examples; hence, we randomize the prompt by manually creating a set of semantically similar instructions. Then, we randomly sample from the instruction set each time.
Step 2: Manual Data Curation. To retain high-quality data for our analysis, we manually filter out generated samples from Step 1 that are invalid or low-quality. Table 1 shows an example where given the original input reasoning question $X_{0}$, the model generated $X_{1}$, where it replaces “Vulcan” with “Neptune”. To study the direct effect by taking the difference of the model output when we provide them $X_{0}$, $R_{0}$ and $X_{1}$, $R_{0}$ as inputs.
Step 3: Generate Reasoning Chain. Finally, to get the indirect effect, we generate the reasoning chain ($R_{0}$, $R_{1}$) for each reasoning problem $X_{0}$ or $X_{1}$ by providing LLMs with some high-level descriptions about each reasoning task and reasoning prompt – “Let’s think step by step”(see App. A.3).
Our study suggests that vanilla LMs (<20B) (in a zero-shot setting) are systematically unfaithful and consistently fail to reason over the mediator (see Table 3). In general, our experiments show a large variation in the causal effects of COT in the final answer depending on the tasks. Models that are instruction tuned or trained on the chain of thought during the pre-training phase have a better indirect effect across different reasoning tasks, suggesting that fine-tuning on CoT can make the model more faithful (see Table 2). Interestingly, similar to Turpin et al. (2023), we observe that there is an inverse scaling for certain tasks. In our case, the indirect effect is worsening with increasingly capable models, indicating that smaller models might be more effective in faithful reasoning.
3 Method
[…]
@@ 3.1 Inference Module
Given a reasoning problem $x\in X$} and instruction prompt $p\in$correct or counterfactual}, our goal is to train models that could generate reasoning steps ($r_{w}$ or $r_{l}$). We propose to adopt Direct Preference Optimization (DPO) (Rafailov et al., 2023), an effective algorithm for aligning language models with implicit rewards. DPO assumes that we only have access to some pairwise preference data $x\rightarrow$$r_{w}$ $>$ $r_{l}$} for each problem $x\in X$. Hence, while training a model ($\pi_{\theta}$) to generate correct reasoning steps, we consider counterfactual and irrelevant reasoning steps as less preferred. Training a DPO model includes two phases: (i) supervised fine-tuning (SFT) and (ii) RL fine-tuning phase.
SFT. We begin by fine-tuning a pre-trained LM with a maximum log-likelihood objective to obtain $\pi_{sft}$.
RL Phase. Contrary to traditional RL approaches, which initially train a reward model and subsequently derive a policy from it, DPO enables extracting policy through implicit reward learning. The preference data of human or artificial annotators is modeled by a learnable implicit reward model $f_{\theta}$ under Bradley-Terry theories (Bradley and Terry, 1952):
During inference, the reasoning module uses the generated reasoning steps by $\pi_{\theta}$ model for a given reasoning problem.
@@ 3.2 Reasoning Module
Given a reasoning question $x\in X$ and reasoning steps $r_{w}$ (correct) and $r_{l}$ (counterfactual) [^3] Please note that in the reasoner module, we only consider counterfactual reasoning steps as negative samples., our goal is to train a model ($\pi_{\gamma}$) that can generate a correct answer $y_{w}$. To encourage our reasoner module to reason faithfully over the reasoning steps, we train the model with a linear combination of two losses: an indirect effect loss and a supervised margin rank loss, $\mathcal{L}=\lambda_{\text{LM}}*\mathcal{L}_{\text{LM}}+\lambda_{\text{counter}}*\mathcal{L}_{\text{counter}}+\lambda_{\text{PREF}}*\mathcal{L}_{\text{PREF}}$ , which we describe below.
@@
where $t$ is the label (indicating which sample in the pair is better), $m$ is the margin, and the indirect effect IE = $h(x,r_{w},y_{w})-h(x,r_{l},y_{w})$ where $h$ is the logits .
@@
Table 2: Causal Effects of generated CoT and reasoning problems on the outputs, with both Natural Indirect Effect (NIE) and Natural Direct Effect (NDE).
@@ 4 Experiments
We conduct the causal mediation analysis on three datasets: StrategyQA (Geva et al., 2021), GSM8K (Cobbe et al., 2021), and Causal Understanding (Suzgun et al., 2023). StrategyQA and Causal Understanding are binary classification tasks. GSM8K dataset contains 8.5K high-quality linguistically diverse grade school math word problems. We evaluate Frodo on four datasets: StrategyQA, Quarel(Tafjord et al., 2018), OpenBookQA (Mihaylov et al., 2018), and QASC (Khot et al., 2020). We report more details about each dataset in App. A.1. For all the datasets, we do not use human-written rationales. We used rationales generated by prior work (Ramnath et al., 2024) using GPT-3 (Text-Davinci-003) for silver rationales for supervision. For counterfactual rationales, we use chain-of-thought prompts on these datasets (refer App. A.3) and sample 2 rationale generations for each training set instance with a temperature of 0.5.
@@
Faithfulness. To evaluate the faithfulness of the rationales generated by the small-sized models, we use LAS (Hase et al., 2020) to measure how well the rationales help a simulator to predict a student’s predictions a’, namely Acc(qr –> a’) - Acc(q –> a’).
@@
We use GPT-4 to generate intervened reasoning problems $X_{1}$ and reasoning chains ($R_{0}$ or $R_{1}$) to perform the causal mediation analysis. In App. A.3, we report the prompt for generating the reasoning chains. Details about the hyperparameter configurations for all the models are in App. A.1.
@@
We perform the causal analysis on a series of language models that are diverse in terms of scale, training, and data: LLaMa-2 (Touvron et al., 2023), Mistral (Jiang et al., 2023), Chat-GPT (Brown et al., 2020a), GPT-4 (OpenAI, 2023), Flan-T5 (Chung et al., 2022), Flan-Alpaca (Chung et al., 2022), Stable-Vicuna (Chiang et al., 2023). We compare Frodo with four strong baselines: (1) SFT + CoT: Finetuning a T5-large or T5-3B on silver rationales, then train another model with LM objective to perform the reasoning, Rainier (Liu et al., 2022), (3) Crystal (Liu et al., 2023), (4) Mario(Ramnath et al., 2024) and (5) Scott (Wang et al., 2023). More details about all the baselines are reported in App. A.2
@@
Table 3: Causal Effects of CoT. The reported results are zero-shot performance. CIE: Controlled Indirect Effect, CDE: Controlled Direct Effect.

### v4

@@ Abstract
Large language models (LLMs) have been shown to perform better when asked to reason step-by-step before answering a question. However, it is unclear to what degree the model’s final answer is faithful to the stated reasoning steps. In this paper, we perform a causal mediation analysis on twelve LLMs to examine how intermediate reasoning steps generated by the LLM influence the final outcome and find that LLMs do not reliably use their intermediate reasoning steps when generating an answer. To address this issue, we introduce Frodo, a framework to tailor small-sized LMs to generate correct reasoning steps and robustly reason over these steps. Frodo consists of an inference module that learns to generate correct reasoning steps using an implicit causal reward function and a reasoning module that learns to faithfully reason over these intermediate inferences using a counterfactual and causal preference objective. Our experiments show that Frodo significantly outperforms four competitive baselines. Furthermore, Frodo improves the robustness and generalization ability of the reasoning LM, yielding higher performance on out-of-distribution test sets. Finally, we find that Frodo’s rationales are more faithful to its final answer predictions than standard supervised fine-tuning.
@@
Debjit Paul, Robert West, Antoine Bosselut, Boi Faltings EPFL {debjit.paul, robert.west, antoine.bosselut, boi.faltings}@epfl.ch
@@ 1 Introduction
Figure 1: An example of our proposed causal analysis to measure the faithfulness of the final output to the CoT generated by the model. We perturbed CoT rationales and studied the causal impact on the model’s behaviour.
Reasoning implicitly involves two steps: identifying the rules and facts (inference chains) necessary to reach a conclusion and then robustly using them to reach said conclusion (Levesque, 1986). Our paper studies whether LLMs reliably use inference chains to arrive at a conclusion. [^1] In our paper, reasoning faithfulness refers to how reliably the model uses its reasoning steps to arrive at a correct answer. In standard CoT, LLMs can generate plausible explanations with the final answer not necessarily guaranteed to follow the reasoning chain or imply a causal relation between the reasoning chain and the model’s outcome (Lyu et al., 2023). Most recent efforts have either focused on the performance of LLMs on various reasoning tasks or their faithfulness in CoT generation, ignoring the sequential relationship between CoT and the final answer (Huang and Chang, 2023; Lanham et al., 2023).
@@
We apply this framework and study the causal impact of CoT rationales on the behaviour of twelve different state-of-the-art LLMs on three different complex reasoning tasks (mathematical, commonsense, and causal understanding). We observe a large variation across tasks and models in how strongly reasoning traces causally affect the model’s prediction. In particular, we find that instruction-tuned models (GPT-3.5-Instruct, Brown et al., 2020b; Mistral-Instruct-7B, Jiang et al., 2023b) have a stronger causal effect on the final answer when conditioned on the reasoning trace than models trained with RLHF (e.g., ChatGPT; Llama-2-7B-Chat, Touvron et al., 2023). Similar to Turpin et al. (2023), when we intervene in the reasoning problem, we observe that ChatGPT and GPT-$3.5$-Instruct are inconsistent at generating plausible reasoning chains. Finally, we find GPT-$4$ (Achiam et al., 2023) only changes its answer $30$% of the time when conditioned on perturbed counterfactual reasoning chains. In Figure 1, we see one example where GPT-4 does not faithfully change its final answer when provided with intervened counterfactual CoT. These results indicate two issues: (i) LLMs can generate unfaithful and implausible reasoning chains, and (ii) LLMs are inconsistent when reasoning over their own generated reasoning traces.
@@
We evaluate Frodo on four reasoning tasks (Quarel, StrategyQA, OpenBookQA, QASC) using multiple model backbones of different scales, and demonstrate that Frodo achieves an absolute accuracy improvement of $2\%\sim 3\%$ over standard supervised fine-tuning or CoT distillation methods. We assess robustness by examining how models alter their answers when intervened with counterfactual reasoning chains. Frodo exhibits significant (+$4.5\%$) improvement in robustness. Finally, Frodo generalizes better to out-of-distribution test sets, showing a +$2.6\%$ performance improvement over supervised fine-tuning. Our code and data are publicly available [^2] https://debjitpaul.github.io/reasoningmatter.
@@ 2 Reasoning Chain as a Mediator
The causal graph is a probabilistic graphical model used to describe how variables interact, expressed by a directed acyclic graph consisting of the sets of nodes ($N$) denoting the variables and a set of directed edges ($E$) indicating the causal relationships between these variables denoting the causality.
@@
It is a method to measure how an independent variable (or treatment) affects a dependent variable (or outcome) mediated by intermediate variables (Pearl, 2001; Robins, 2003). Causal mediation analysis aims to decompose the total effect of the independent variable ($X$) on the dependent variable ($Y$) into two components: the direct effect and the indirect effect (Pearl, 2001). In this work, we view the reasoning process as a causal graph, framing the input (reasoning problem) $X$ and the output $Y$ as random variables and the reasoning steps as mediator variable $R$. We use mediation analysis to interpret the role of reasoning steps as mediators between model inputs and model outputs. Let $X_{0}$ denote the initial reasoning problem, $R_{0}$ the reasoning chain given $X_{0}$. Let $Y_{00}$ denote the potential outcome when the treatment and mediator variables are $X_{0}$ and $R_{0}$, respectively. Meanwhile, $Y_{01}$ denotes the potential outcome when treatment is set to $X_{0}$, and $R_{1}$ is the reasoning chain for the reasoning problem $X_{1}$.
@@
Figure 2: Causal graph for natural language reasoning, modeling $P(Y|do(x))$. $X_{0}$: original reasoning problem, $X_{1}$: intervened reasoning problem. $R_{0}$: Reasoning steps for $X_{0}$, $R_{1}$: Reasoning steps for $X_{1}$. Outputs $Y_{00}$ or $Y_{01}$ are model outputs given $X_{0}$ and $R_{0}$ or $X_{0}$ and $R_{1}$.
@@
Direct Effect (DE) measures how much an intervention $X$ changes an outcome variable $Y$ directly, without passing through a hypothesized mediator $R$. The direct effect of $X=X_{0}$ on $Y$ can be defined as $\mathbb{E}[Y_{00}-Y_{10}]$, which can be seen as the correctness comparison between the two potential outcomes given two different treatments, i.e., $X=X_{0}$ and $X=X_{1}$. It is computed by applying the intervention $X$ but holding $R$ fixed to its original value ($R_{0}$).
@@
Indirect Effect (IE) measures how much an intervention $X$ changes $Y$ indirectly through $R$. The indirect effect can be defined as IE = $\mathbb{E}[Y_{00}-Y_{01}]$. It is computed by setting $R$ to its value under the intervention $X$ while keeping everything else to its original value.
@@
More concretely, according to Pearl (2001), in our scenario, a high direct effect means the model emphasizes the reasoning problem more than the reasoning steps. In contrast, a high indirect effect means the model emphasizes the reasoning steps more than the problem input.
Reasoning Intervention.
Following Pearl (2001), we conduct counterfactual reasoning to measure the causal effect of a treatment variable on a response variable. We first perform targeted interventions on the input text $X$ and measure their effect on the reasoning outcome $Y$ by keeping $R$ fixed (direct effect). Further, we also perform interventions on the mediator $R$ and measure their effect on $Y$ (indirect effect). We perform the following steps to automatically generate an intervention on $X$ and $R$.
Step 1: Intervention Data Generation. We use a large language model (GPT-4) to automatically generate an alternative value $X_{1}$ for the treatment variable. [^3] See Table. 23 for details on task-specific interventions. The input to LLM includes instruction and few-shot examples, taking the format shown in Table 20. LLMs can be sensitive to instructions and few-shot examples; hence, we randomize the prompt by manually creating a set of semantically similar instructions. Then, we randomly sample from the instruction set each time.
Step 2: Manual Data Curation. To retain high-quality data for our analysis, we manually filter out generated samples from Step 1 that are invalid or low-quality. Table 1 shows an example where given the original input reasoning question $X_{0}$, the model generated $X_{1}$, where it replaces “Vulcan” with “Neptune”.
Step 3: Generate Reasoning Chain. Finally, to get the indirect effect, we generate the reasoning chain ($R_{0}$, $R_{1}$) for each reasoning problem $X_{0}$ or $X_{1}$ by providing LLMs with some high-level descriptions about each reasoning task and reasoning prompt – “Let’s think step by step”(see App. Table. 19).
@@
Table 1: An example from StrategyQA dataset, where $X_{1}$ = intervened reasoning problem; $R_{0}$ and $R_{1}$ = reasoning steps (generated by GPT-4).
Our study suggests that vanilla LMs (<20B) (in a zero-shot setting) are systematically unfaithful and consistently fail to reason over the mediator (see Table 3). In general, our experiments show a large variation in the causal effects of COT in the final answer depending on the tasks. Models that are instruction tuned or trained on the chain of thought during the pre-training phase have a better indirect effect across different reasoning tasks, suggesting that fine-tuning on CoT can make the model more faithful (see Table 2). Interestingly, similar to Turpin et al. (2023), we observe an inverse scaling for certain tasks. In our case, the indirect effect is worsening with increasingly capable models, indicating that smaller models might be more effective in faithful reasoning.
3 FRODO
In this section, we introduce Frodo, a framework that tailors small-sized LMs (<10B parameters) to be strong rationalizers and perform reasoning faithfully over the rationales. Frodo aims to improve the synergy between the reasoning chain and the final answer. We first describe how we obtain silver reasoning chains from LLMs (§$\S$3.1). Then, we introduce our inference module that trains a model to generate rationales (§$\S$3.1) followed by the reasoner module and its training objectives (§$\S$3.2).
@@ 3.1 Inference Module
Given a reasoning problem $x\in X$} and instruction prompt $p\in$correct or counterfactual}, our goal is to train models that could generate reasoning steps ($r_{w}$ or $r_{l}$). We propose to adopt Direct Preference Optimization (DPO) (Rafailov et al., 2023), an effective algorithm for aligning language models with implicit rewards. DPO assumes that we only have access to some pairwise preference data $x\rightarrow$$r_{w}$ $>$ $r_{l}$} for each problem $x\in X$. Hence, while training a model ($\pi_{\theta}$) to generate correct reasoning steps, we consider counterfactual and irrelevant reasoning steps as less preferred. Training a DPO model includes two phases: (i) supervised fine-tuning (SFT) and (ii) Preference Learning (PL) phase.
SFT.
We begin by fine-tuning a pre-trained LM with a maximum log-likelihood objective to obtain $\pi_{sft}$.
PL Phase.
Contrary to traditional RL approaches, which initially train a reward model and subsequently derive a policy from it, DPO enables extracting policy through implicit reward learning. DPO adopts a binary classification loss:
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{\{x,r_{w}>r_{l}\}}\text{log}\sigma(f_{\theta}(r_{w},x)-f_{\theta}(r_{l},x))$ (1)
where $f_{\theta}$ is the implicit reward model. Intuitively, the gradient of the loss function $\mathcal{L}_{\text{DPO}}$ increases the likelihood of the preferred completions $r_{w}$ and decreases the likelihood of counterfactual reasoning chains $r_{l}$. See Appendix A.1 for more details. During inference, the reasoning module uses the generated reasoning steps by $\pi_{\theta}$ model for a given reasoning problem.
@@ 3.2 Reasoning Module
Given a reasoning question $x\in X$ and reasoning steps $r_{w}$ (correct) and $r_{l}$ (counterfactual) [^4] Please note that in the reasoner module, we only consider counterfactual reasoning steps as negative samples., our goal is to train a model ($\pi_{\gamma}$) that can generate a correct answer $y_{w}$. To encourage our reasoner module to reason faithfully over the reasoning steps, we train the model with a linear combination of three losses: an indirect effect loss and a supervised margin rank loss, $\mathcal{L}=\lambda_{\text{LM}}*\mathcal{L}_{\text{LM}}+\lambda_{\text{counter}}*\mathcal{L}_{\text{counter}}+\lambda_{\text{PREF}}*\mathcal{L}_{\text{PREF}}$ , which we describe below.
@@
where $t$ is the label (indicating which sample in the pair is better)=$1$, $m$ is the margin=$1.0$, and the indirect effect IE = $h(x,r_{w},y_{w})-h(x,r_{l},y_{w})$ where $h$ is the logits.
@@
Table 2: Causal Effects of generated CoT and reasoning problems on the outputs, with both Natural Indirect Effect (NIE) and Natural Direct Effect (NDE). COT (%) represents the accuracy of the models.
@@ 4 Experiments
We conduct the causal mediation analysis on three datasets: StrategyQA (Geva et al., 2021), GSM8K (Cobbe et al., 2021), and Causal Understanding (Suzgun et al., 2023). We evaluate Frodo on four datasets: StrategyQA, Quarel(Tafjord et al., 2018), OpenBookQA (Mihaylov et al., 2018), and QASC (Khot et al., 2020). We report more details about each dataset in App. A.3. For all the datasets, we do not use human-written rationales. We used rationales generated by prior work (Ramnath et al., 2024) using GPT-3 (Text-Davinci-003) as silver rationales for supervision. For counterfactual rationales, we use chain-of-thought prompts on these datasets (Table. 19) and sample 2 rationales for each training instance with a temperature of 0.5.
@@
We use GPT-4 to generate intervened reasoning problems $X_{1}$ and reasoning chains ($R_{0}$ or $R_{1}$) to perform the causal mediation analysis. We report the prompts used in Table.19, 20 and hyperparameters in App. A.3.
@@
We perform the causal analysis on a series of language models that are diverse in terms of scale, training, and data: LLaMa-2 (Touvron et al., 2023), Mistral (Jiang et al., 2023b), Chat-GPT (Brown et al., 2020a), GPT-4 (OpenAI, 2023), Flan-T5 (Chung et al., 2022), Flan-Alpaca (Chung et al., 2022), Stable-Vicuna (Chiang et al., 2023). We compare Frodo with four strong baselines: (1) SFT + CoT: Finetuning a T5-large or T5-3B or LLaMa-2-7B with LoRA or Mistral-7B with LoRA on silver rationales, then train another model with LM objective to perform the reasoning, (2) Rainier (Liu et al., 2022), where they used PPO (Proximal Policy Optimization) inference modules, and for the reasoning module, they used SFT (simple finetuning). (3) Crystal (Liu et al., 2023) used PPO to train both inference and reasoning modules, (4) Mario (Ramnath et al., 2024) used QUARK, a multi-reward reinforcement learning method, and for the reasoning module, they used SFT and (5) SCOTT(Wang et al., 2023) used simple-finetuning with contrastive decoding. More details about all the baselines are reported in App. A.4.
@@
Table 3: Causal Effects of CoT. The reported results are zero-shot performance. CIE: Controlled Indirect Effect, CDE: Controlled Direct Effect. The p-value represents the significance of the results